# FRISM: Fine-Grained Reasoning Injection via Subspace-Level Model Merging for Vision–Language Models

### v1

@@ Appendix E Analysis of IP-Merging’s Failure Cases
Appendix F Experimental results on Llama-based models
@@
Results. The results are presented in Tab. 6. FRISM shows significant improvements in VL reasoning tasks compared to existing methods while still reserving good VL perception capabilities. These results further demonstrate the applicability of FRISM to different kinds of models.
Figure 7: Keyword Analysis of different merging methods for (a) reasoning-related keywords and (b) vision-related keywords.
Appendix G Theoretical Analysis
G.1 Subspace-level filtering mechanism
@@
$\begin{split}&\mathcal{L}_{\text{vis}}({\theta}_{\text{vlm}}+\Delta{\theta})\\ \approx&\mathcal{L}_{\text{vis}}({\theta}_{\text{vlm}})+\nabla\mathcal{L}_{\text{vis}}({\theta}_{\text{vlm}})^{\top}\Delta\theta+\frac{1}{2}\Delta{\theta}^{\top}\mathbf{H}\Delta{\theta}\\ \approx&\frac{1}{2}\Delta{\theta}^{\top}\mathbf{H}\Delta{\theta},\end{split}$ (16)
@@
$\delta_{\text{vis}}\left(\Delta\theta\right)=\frac{1}{2}\Delta{\theta}^{\top}\mathbf{H}\Delta{\theta}$ (17)
@@
$\Delta\theta=\sum_{i=1}^{N}\lambda_{i}B_{i},$ (18)
@@
$\mathcal{L}=\mathcal{L}_{\text{vis}}-\alpha\mathcal{L}_{\text{reason}}=\frac{1}{2}\Delta\theta^{\top}\mathbf{H}\Delta\theta-\alpha\|\Delta\theta\|_{F}^{2}.$ (19)
@@
$\mathcal{L}=\frac{1}{2}\sum_{i,j}\lambda_{i}\lambda_{j}\text{Tr}(B_{i}^{\top}\mathbf{H}B_{j})-\alpha\sum_{i,j}\lambda_{i}\lambda_{j}\text{Tr}(B_{i}^{\top}B_{j}).$ (20)
@@
$\mathcal{L}\approx\sum_{i=1}^{N}\left(\frac{1}{2}\lambda_{i}^{2}h_{i}-\alpha\lambda_{i}^{2}\|B_{i}\|_{F}^{2}\right),$ (22)
@@
Analyzing the gradient of Eq. 22 with respect to the gate $\lambda_{i}$:
$\frac{\partial\mathcal{L}}{\partial\lambda_{i}}=\left(h_{i}-2\alpha\|B_{i}\|_{F}^{2}\right)\cdot\lambda_{i}.$ (23)
@@
G.2 Loss term analysis
@@
$\mathcal{L}_{\text{inject}}=-\sum_{l=1}^{L}\|\mathbf{S}_{\text{eff}}^{(l)}\|^{2}=-\sum_{l=1}^{L}\|\mathrm{Sigmoid}(\mathbf{g}^{(l)})\odot\mathbf{S}^{(l)}\|^{2},$ (24)
@@
$\theta_{\text{merged}}=\theta_{\text{base}}+\tau_{\text{vlm}}+\Delta\theta_{\text{lrm}}^{\text{eff}}$ (25)
@@
$\begin{split}J(\lambda)&=\|\theta_{\text{lrm}}-\theta_{\text{merged}}\|_{F}^{2}\\ &=\|\tau_{\text{lrm}}-(\tau_{\text{vlm}}+\Delta\theta_{\text{lrm}}^{\text{eff}})\|_{F}^{2}\end{split}$ (26)
@@
$\begin{split}J(\lambda)&=\|\Delta_{\text{res}}-\tau_{\text{vlm}}\|_{F}^{2}\\ &=\|\Delta_{\text{res}}\|_{F}^{2}+\|\tau_{\text{vlm}}\|_{F}^{2}-2\langle\Delta_{\text{res}},\tau_{\text{vlm}}\rangle\end{split}$ (27)
@@
$\begin{split}&\langle B_{i},\tau_{\text{vlm}}\rangle\approx 0,\\ &\langle\Delta_{\text{res}},\tau_{\text{vlm}}\rangle\approx 0.\end{split}$ (28)
@@
$J(\lambda)\approx\sum_{i=1}^{N}(1-\lambda_{i})^{2}\|B_{i}\|_{F}^{2}+C$ (29)
@@
Figure 8: Qualitative comparison of attention maps across different model merging methods on a sample from MathVista dataset.
Appendix H Visualization Results
In Fig. 8, we present a qualitative comparison of attention maps across different methods on a sample from the MathVista (Lu et al., 2024a) dataset. We visualize the regions focused on by the models in the image. As observed, Base VLM focuses on the markings of the measuring cup. The models merged by the baseline methods usually exhibit a shift in the visual attention region. In contrast, our FRISM demonstrates visual attention regions similar to those of the base model. Furthermore, in terms of text attention regions, baseline methods tend to exhibit attention dispersion and focus on the format instrctions or syntactic artifacts, ignoring the critical chain-of-thought triggers. In contrast, FRISM maintains high attention weights on these key reasoning keywords, including “reason” and “step”. This demonstrates that our method achieves effective reasoning capability injection without compromising visual robustness.
Appendix I FRISM Technique Details
I.1 Implementation details
@@
$\mathcal{L}_{\text{inject}}^{\text{init}}=-\sum_{l=1}^{L}\|\mathrm{Sigmoid}(\mathbf{g_{\text{init}}}^{(l)})\odot\mathbf{S}^{(l)}\|^{2}.$ (30)
@@
$\mathcal{L}_{\text{inject}}^{\text{init}}=-\sum_{l=1}^{L}\|\frac{1}{2}\mathbf{S}^{(l)}\|^{2}.$ (31)
@@
$\mathcal{L}_{\text{inject}}^{\text{training}}=-\frac{\sum_{l=1}^{L}\|\mathrm{Sigmoid}(\mathbf{g}^{(l)})\odot\mathbf{S}^{(l)}\|^{2}}{\mathcal{L}_{\text{inject}}^{\text{init}}}.$ (32)
I.2 Hyper-parameter analysis
@@
Injection Coefficient $\alpha$ balances the dual objectives of visual preservation ($\mathcal{L}_{\text{distill}}$) and reasoning maximization ($\mathcal{L}_{\text{inject}}$). Our theoretical analysis in Appendix G, Eq 23 reveals its physical significance as a threshold controller for the subspace filter. $\frac{\partial\mathcal{L}}{\partial\lambda_{i}}=\left(h_{i}-2\alpha\|B_{i}\|_{F}^{2}\right)\cdot\lambda_{i}$.
@@
If $\lambda_{\text{lrm}}$ is set too low, even with $g\rightarrow\infty$, the effective singular value $S_{\text{eff}}$ may remain too small to influence the model’s behavior, leading to under-fitting of the reasoning task. Conversely, an excessively high $\lambda_{\text{lrm}}$ can lead to unstable initialization before the gates adapt. In our experiments, we usually set $\lambda_{\text{lrm}}$ close to $0.2$, which is adjustable depending on the magnitude of the LRM task vector $\tau_{\text{lrm}}$.
Appendix J More results for keyword analysis
Table 7: Count of Keywords generated by different methods when merging Qwen2.5-VL-32B with QwQ-32B models.
@@
In Tab.7, we show more keyword analysis results on MathVision when merging Qwen2.5-VL-32B-Instruct and QwQ-32B. Our FRISM shows both strong self-reflection and visual focus capabilities. A visualized comparison is shown in Fig. 7.
Appendix K Case Study
We present a case study from the MathVision (Wang et al., 2024a) dataset. The image in the question, prompt for the model, and the correct answer are shown in Box K. In this section, we compare the responses generated by the base VLM Qwen2.5-VL-7B-Instruct and the models merged with DeepSeek-R1-Distill-Qwen-7B through Task Arithmetic, IP-Merging, and ours. The generated responses are respectively shown in Box K, K, K, K. It can be seen that the response generated by ours shows reasoning and re-thinking capabilities and finally outputs the correct answer.
@@
Box K.2: Response generated by base VLM (Qwen2.5-VL-7B-Instruct). To solve for the measure of angle $\angle XAY$, we analyze the given geometric configuration. 1. Identify the given information: The circle has radius $r$. Points $X$ and $Y$ lie on the circumference of the circle. Point AA also lies on the circumference of the circle. $XY=r$. $XY\perp AY$. 2. Analyze the right triangle: Since $XY\perp AY$, triangle AXYAXY is a right triangle with $\angle AXY=90^{\circ}.$ 3. Determine the relationship between the sides: In a right triangle, the hypotenuse is the longest side. Here, AYAY is the hypotenuse since it connects two points on the circumference and is perpendicular to $XY$. 4. Use the properties of the circle: Because $X$ and $Y$ lie on the circumference and $XY=r$, the segment XYXY is a chord of the circle. Given that $XY\perp AY$, point AA lies directly above or below the midpoint of XYXY along a diameter of the circle. 5. Calculate the angle: In a right triangle inscribed in a semicircle, the angle subtended by the diameter at any point on the circumference is always $90^{\circ}$. Therefore, $\angle XAY=90^{\circ}.$ $\boxed{90^{\circ}}$
@@
Box K.3: Response generated by Task Arithmetic. To solve for the angle $\angle XAY$ in the given problem, we analyze the geometric configuration and apply properties of circles and right triangles. 1. Identify the given information: The circle has radius $r$. Points $X$, $Y$, and AA lie on the circumference of the circle. $XY=r$. $XY\perp AY$. 2. Analyze triangle $\triangle AXY$: Since XYXY is a chord of the circle and $XY\perp AY$, segment AYAY must be a radius of the circle because the perpendicular from the center of a circle to a chord bisects the chord. Therefore, $AY=r.$ 3. Determine the type of triangle: Triangle $\triangle AXY$ is a right triangle with $\angle AXY=90^{\circ}$ because $XY\perp AY$. Since both AXAX and AYAY are radii of the circle, we have $AX=AY=r.$ 4. Calculate the angles in the triangle: Triangle $\triangle AXY$ is isosceles, so the base angles $\angle XAY$ and $\angle AYX$ are equal. Let $\angle XAY=\theta.$ Then, $\angle AYX=\theta.$ 5. Use the angle sum of a triangle: The sum of the angles in a triangle is $180^{\circ}$. Hence, $\theta+\theta+90^{\circ}=180^{\circ}.$ Simplifying, $2\theta+90^{\circ}=180^{\circ},$ which gives $2\theta=90^{\circ}\quad\Rightarrow\quad\theta=45^{\circ}.$ Thus, the angle $\angle XAY$ is $\boxed{45^{\circ}}.$
@@
Box K.4: Response generated by the model obtained through IP-Merging. To solve for the measure of angle $\angle XAY$, we analyze the given geometric configuration. 1. Identify the given information: The circle has a radius $r$. Points $X$ and $Y$ lie on the circumference of the circle. Point AA is also on the circumference of the circle. $XY=r$. $XY\perp AY$. 2. Analyze the right triangle: Since $XY\perp AY$, triangle AXYAXY is a right triangle with $\angle AXY=90^{\circ}.$ 3. Determine the relationship between the sides: In a right triangle, the hypotenuse is the longest side. Here, AYAY is the hypotenuse because it connects two points on the circumference of the circle and is perpendicular to $XY$. 4. Use the properties of the circle: Since $X$ and $Y$ are points on the circumference and $XY=r$, the segment XYXY is a chord of the circle. Given that $XY\perp AY$, point AA must lie directly above or below the midpoint of XYXY along a diameter of the circle. 5. Calculate the angle: In a right triangle inscribed in a semicircle, the angle subtended by the diameter at any point on the circumference is always $90^{\circ}$. Therefore, $\angle XAY=90^{\circ}.$ Thus, the measure of angle $\angle XAY$ is $\boxed{90^{\circ}}.$
@@
Box K.5: Response generated by the model obtained through FRISM. To solve the problem, we analyze the given geometric configuration using properties of circles and right triangles. 1. Identify the given information: The circle has radius $r$. Points $X$, $Y$, and AA lie on the circumference of the circle. $XY=r$. $XY\perp AY$. 2. Analyze triangle $\triangle XYO$: Since $XY=r$ and $XO=YO=r$ (both are radii of the circle), triangle $\triangle XYO$ is equilateral. 3. Determine the central angle subtended by arc XYXY: In an equilateral triangle inscribed in a circle, each side subtends a central angle of $60^{\circ}$. Therefore, $\angle XOY=60^{\circ}.$ 4. Analyze the right triangle $\triangle AXY$: Since $XY\perp AY$, triangle $\triangle AXY$ is a right triangle with $\angle AYX=90^{\circ}.$ Let $\angle XAY=\theta.$ 5. Use the angle sum of a triangle: The three angles of $\triangle AXY$ are $\angle AYX=90^{\circ},\quad\angle XAY=\theta,\quad\angle AXY=90^{\circ}-\theta.$ Their sum is $90^{\circ}+\theta+(90^{\circ}-\theta)=180^{\circ},$ which confirms the setup is consistent. 6. Determine $\angle XAY$: Since AA lies on the circumference, the inscribed angle $\angle XAY$ subtends the same arc XYXY as the central angle $\angle XOY$. An inscribed angle equals half the corresponding central angle, so $\angle XAY=\frac{1}{2}\angle XOY=\frac{60^{\circ}}{2}=30^{\circ}.$ $\boxed{\angle XAY=30^{\circ}}$
@@
Despite the convincing results, the proposed FRISM suffers from several limitations. On the one hand, though negligible compared to post-training methods, FRISM requires some training costs. On the other hand, as a common limitation of task vector-based merging methods, FRISM cannot be generalized to models with different sizes or structures. Further transferring reasoning capabilities from reasoning language models to models with more modalities, model sizes, or even different structures is a significant future direction.

### v2

@@ D.1 More experiments comparing with more models on more benchmarks
D.2 Experimental results on Llama-based models
@@
Results. The results are presented in Tab. A4. FRISM shows significant improvements in VL reasoning tasks compared to existing methods while still reserving good VL perception capabilities. These results further demonstrate the applicability of FRISM to different kinds of models.
D.3 More comparisons between subspace-level and layer-wise merging
Following the configuration of Fig. 3, we further plot the results by merging the VLMs and the task vectors from another model, OpenThinker3-7B (Guha et al., 2025), in Fig. A1. The results further demonstrate the effectiveness of merging in the subspace level.
Figure A1: Impact of scaling coefficients across subspace ranks when merging Qwen2.5-VL-7B-Instruct and the subspaces of OpenThinker task vector.
Appendix E Analysis of IP-Merging’s Failure Cases
$\lambda_{n}=\frac{\sum_{i=1}^{d}\sigma_{\text{vlm},i}^{n}}{\sum_{i=1}^{d}\sigma_{\text{lrm},i}^{n}},$ (A2)
Figure A2: Keyword Analysis of different merging methods for (a) reasoning-related keywords and (b) vision-related keywords.
Figure A3: Qualitative comparison of attention maps across different model merging methods on a sample from MathVista dataset.
Appendix F Theoretical Analysis
F.1 Subspace-level filtering mechanism
@@
$\begin{split}&\mathcal{L}_{\text{vis}}({\theta}_{\text{vlm}}+\Delta{\theta})\\ \approx&\mathcal{L}_{\text{vis}}({\theta}_{\text{vlm}})+\nabla\mathcal{L}_{\text{vis}}({\theta}_{\text{vlm}})^{\top}\Delta\theta+\frac{1}{2}\Delta{\theta}^{\top}\mathbf{H}\Delta{\theta}\\ \approx&\frac{1}{2}\Delta{\theta}^{\top}\mathbf{H}\Delta{\theta},\end{split}$ (A3)
@@
$\delta_{\text{vis}}\left(\Delta\theta\right)=\frac{1}{2}\Delta{\theta}^{\top}\mathbf{H}\Delta{\theta}$ (A4)
@@
$\Delta\theta=\sum_{i=1}^{N}\lambda_{i}B_{i},$ (A5)
@@
$\mathcal{L}=\mathcal{L}_{\text{vis}}-\alpha\mathcal{L}_{\text{reason}}=\frac{1}{2}\Delta\theta^{\top}\mathbf{H}\Delta\theta-\alpha\|\Delta\theta\|_{F}^{2}.$ (A6)
@@
$\mathcal{L}=\frac{1}{2}\sum_{i,j}\lambda_{i}\lambda_{j}\text{Tr}(B_{i}^{\top}\mathbf{H}B_{j})-\alpha\sum_{i,j}\lambda_{i}\lambda_{j}\text{Tr}(B_{i}^{\top}B_{j}).$ (A7)
@@
$\mathcal{L}\approx\sum_{i=1}^{N}\left(\frac{1}{2}\lambda_{i}^{2}h_{i}-\alpha\lambda_{i}^{2}\|B_{i}\|_{F}^{2}\right),$ (A9)
@@
Analyzing the gradient of Eq. A9 with respect to the gate $\lambda_{i}$:
$\frac{\partial\mathcal{L}}{\partial\lambda_{i}}=\left(h_{i}-2\alpha\|B_{i}\|_{F}^{2}\right)\cdot\lambda_{i}.$ (A10)
@@
F.2 Loss term analysis
@@
$\mathcal{L}_{\text{inject}}=-\sum_{l=1}^{L}\|\mathbf{S}_{\text{eff}}^{(l)}\|^{2}=-\sum_{l=1}^{L}\|\mathrm{Sigmoid}(\mathbf{g}^{(l)})\odot\mathbf{S}^{(l)}\|^{2},$ (A11)
@@
$\theta_{\text{merged}}=\theta_{\text{base}}+\tau_{\text{vlm}}+\Delta\theta_{\text{lrm}}^{\text{eff}}$ (A12)
@@
$\begin{split}J(\lambda)&=\|\theta_{\text{lrm}}-\theta_{\text{merged}}\|_{F}^{2}\\ &=\|\tau_{\text{lrm}}-(\tau_{\text{vlm}}+\Delta\theta_{\text{lrm}}^{\text{eff}})\|_{F}^{2}\end{split}$ (A13)
@@
$\begin{split}J(\lambda)&=\|\Delta_{\text{res}}-\tau_{\text{vlm}}\|_{F}^{2}\\ &=\|\Delta_{\text{res}}\|_{F}^{2}+\|\tau_{\text{vlm}}\|_{F}^{2}-2\langle\Delta_{\text{res}},\tau_{\text{vlm}}\rangle\end{split}$ (A14)
@@
$\begin{split}&\langle B_{i},\tau_{\text{vlm}}\rangle\approx 0,\\ &\langle\Delta_{\text{res}},\tau_{\text{vlm}}\rangle\approx 0.\end{split}$ (A15)
@@
$J(\lambda)\approx\sum_{i=1}^{N}(1-\lambda_{i})^{2}\|B_{i}\|_{F}^{2}+C$ (A16)
@@
Appendix G Visualization Results
In Fig. A3, we present a qualitative comparison of attention maps across different methods on a sample from the MathVista (Lu et al., 2024a) dataset. We visualize the regions focused on by the models in the image. As observed, Base VLM focuses on the markings of the measuring cup. The models merged by the baseline methods usually exhibit a shift in the visual attention region. In contrast, our FRISM demonstrates visual attention regions similar to those of the base model. Furthermore, in terms of text attention regions, baseline methods tend to exhibit attention dispersion and focus on the format instrctions or syntactic artifacts, ignoring the critical chain-of-thought triggers. In contrast, FRISM maintains high attention weights on these key reasoning keywords, including “reason” and “step”. This demonstrates that our method achieves effective reasoning capability injection without compromising visual robustness.
Appendix H FRISM Technique Details
H.1 Implementation details
@@
$\mathcal{L}_{\text{inject}}^{\text{init}}=-\sum_{l=1}^{L}\|\mathrm{Sigmoid}(\mathbf{g_{\text{init}}}^{(l)})\odot\mathbf{S}^{(l)}\|^{2}.$ (A17)
@@
$\mathcal{L}_{\text{inject}}^{\text{init}}=-\sum_{l=1}^{L}\|\frac{1}{2}\mathbf{S}^{(l)}\|^{2}.$ (A18)
@@
$\mathcal{L}_{\text{inject}}^{\text{training}}=-\frac{\sum_{l=1}^{L}\|\mathrm{Sigmoid}(\mathbf{g}^{(l)})\odot\mathbf{S}^{(l)}\|^{2}}{\mathcal{L}_{\text{inject}}^{\text{init}}}.$ (A19)
H.2 Hyper-parameter analysis
@@
Injection Coefficient $\alpha$ balances the dual objectives of visual preservation ($\mathcal{L}_{\text{distill}}$) and reasoning maximization ($\mathcal{L}_{\text{inject}}$). Our theoretical analysis in Appendix F, Eq A10 reveals its physical significance as a threshold controller for the subspace filter. $\frac{\partial\mathcal{L}}{\partial\lambda_{i}}=\left(h_{i}-2\alpha\|B_{i}\|_{F}^{2}\right)\cdot\lambda_{i}$.
@@
If $\lambda_{\text{lrm}}$ is set too low, even with $g\rightarrow\infty$, the effective singular value $S_{\text{eff}}$ may remain too small to influence the model’s behavior, leading to under-fitting of the reasoning task. Conversely, an excessively high $\lambda_{\text{lrm}}$ can lead to unstable initialization before the gates adapt. In our experiments, the global merging coefficient $\lambda_{\text{lrm}}$ is searched in $\{0.1,0.15,0.2\}$. which is adjustable depending on the magnitude of the LRM task vector $\tau_{\text{lrm}}$.
Appendix I More results for keyword analysis
Table A5: Count of Keywords generated by different methods when merging Qwen2.5-VL-32B with QwQ-32B models.
@@ Appendix I More results for keyword analysis
In Tab.A5, we show more keyword analysis results on MathVision when merging Qwen2.5-VL-32B-Instruct and QwQ-32B. Our FRISM shows both strong self-reflection and visual focus capabilities. A visualized comparison is shown in Fig. A2.
Appendix J Case Study
We present a case study from the MathVision (Wang et al., 2024a) dataset. The image in the question, prompt for the model, and the correct answer are shown in Box J. In this section, we compare the responses generated by the base VLM Qwen2.5-VL-7B-Instruct and the models merged with DeepSeek-R1-Distill-Qwen-7B through Task Arithmetic, IP-Merging, and ours. The generated responses are respectively shown in Box J, J, J, J. It can be seen that the response generated by ours shows reasoning and re-thinking capabilities and finally outputs the correct answer.
@@
Box J.2: Response generated by base VLM (Qwen2.5-VL-7B-Instruct). To solve for the measure of angle $\angle XAY$, we analyze the given geometric configuration. 1. Identify the given information: The circle has radius $r$. Points $X$ and $Y$ lie on the circumference of the circle. Point AA also lies on the circumference of the circle. $XY=r$. $XY\perp AY$. 2. Analyze the right triangle: Since $XY\perp AY$, triangle AXYAXY is a right triangle with $\angle AXY=90^{\circ}.$ 3. Determine the relationship between the sides: In a right triangle, the hypotenuse is the longest side. Here, AYAY is the hypotenuse since it connects two points on the circumference and is perpendicular to $XY$. 4. Use the properties of the circle: Because $X$ and $Y$ lie on the circumference and $XY=r$, the segment XYXY is a chord of the circle. Given that $XY\perp AY$, point AA lies directly above or below the midpoint of XYXY along a diameter of the circle. 5. Calculate the angle: In a right triangle inscribed in a semicircle, the angle subtended by the diameter at any point on the circumference is always $90^{\circ}$. Therefore, $\angle XAY=90^{\circ}.$ $\boxed{90^{\circ}}$
@@
Box J.3: Response generated by Task Arithmetic. To solve for the angle $\angle XAY$ in the given problem, we analyze the geometric configuration and apply properties of circles and right triangles. 1. Identify the given information: The circle has radius $r$. Points $X$, $Y$, and AA lie on the circumference of the circle. $XY=r$. $XY\perp AY$. 2. Analyze triangle $\triangle AXY$: Since XYXY is a chord of the circle and $XY\perp AY$, segment AYAY must be a radius of the circle because the perpendicular from the center of a circle to a chord bisects the chord. Therefore, $AY=r.$ 3. Determine the type of triangle: Triangle $\triangle AXY$ is a right triangle with $\angle AXY=90^{\circ}$ because $XY\perp AY$. Since both AXAX and AYAY are radii of the circle, we have $AX=AY=r.$ 4. Calculate the angles in the triangle: Triangle $\triangle AXY$ is isosceles, so the base angles $\angle XAY$ and $\angle AYX$ are equal. Let $\angle XAY=\theta.$ Then, $\angle AYX=\theta.$ 5. Use the angle sum of a triangle: The sum of the angles in a triangle is $180^{\circ}$. Hence, $\theta+\theta+90^{\circ}=180^{\circ}.$ Simplifying, $2\theta+90^{\circ}=180^{\circ},$ which gives $2\theta=90^{\circ}\quad\Rightarrow\quad\theta=45^{\circ}.$ Thus, the angle $\angle XAY$ is $\boxed{45^{\circ}}.$
@@
Box J.4: Response generated by the model obtained through IP-Merging. To solve for the measure of angle $\angle XAY$, we analyze the given geometric configuration. 1. Identify the given information: The circle has a radius $r$. Points $X$ and $Y$ lie on the circumference of the circle. Point AA is also on the circumference of the circle. $XY=r$. $XY\perp AY$. 2. Analyze the right triangle: Since $XY\perp AY$, triangle AXYAXY is a right triangle with $\angle AXY=90^{\circ}.$ 3. Determine the relationship between the sides: In a right triangle, the hypotenuse is the longest side. Here, AYAY is the hypotenuse because it connects two points on the circumference of the circle and is perpendicular to $XY$. 4. Use the properties of the circle: Since $X$ and $Y$ are points on the circumference and $XY=r$, the segment XYXY is a chord of the circle. Given that $XY\perp AY$, point AA must lie directly above or below the midpoint of XYXY along a diameter of the circle. 5. Calculate the angle: In a right triangle inscribed in a semicircle, the angle subtended by the diameter at any point on the circumference is always $90^{\circ}$. Therefore, $\angle XAY=90^{\circ}.$ Thus, the measure of angle $\angle XAY$ is $\boxed{90^{\circ}}.$
@@
Box J.5: Response generated by the model obtained through FRISM. To solve the problem, we analyze the given geometric configuration using properties of circles and right triangles. 1. Identify the given information: The circle has radius $r$. Points $X$, $Y$, and AA lie on the circumference of the circle. $XY=r$. $XY\perp AY$. 2. Analyze triangle $\triangle XYO$: Since $XY=r$ and $XO=YO=r$ (both are radii of the circle), triangle $\triangle XYO$ is equilateral. 3. Determine the central angle subtended by arc XYXY: In an equilateral triangle inscribed in a circle, each side subtends a central angle of $60^{\circ}$. Therefore, $\angle XOY=60^{\circ}.$ 4. Analyze the right triangle $\triangle AXY$: Since $XY\perp AY$, triangle $\triangle AXY$ is a right triangle with $\angle AYX=90^{\circ}.$ Let $\angle XAY=\theta.$ 5. Use the angle sum of a triangle: The three angles of $\triangle AXY$ are $\angle AYX=90^{\circ},\quad\angle XAY=\theta,\quad\angle AXY=90^{\circ}-\theta.$ Their sum is $90^{\circ}+\theta+(90^{\circ}-\theta)=180^{\circ},$ which confirms the setup is consistent. 6. Determine $\angle XAY$: Since AA lies on the circumference, the inscribed angle $\angle XAY$ subtends the same arc XYXY as the central angle $\angle XOY$. An inscribed angle equals half the corresponding central angle, so $\angle XAY=\frac{1}{2}\angle XOY=\frac{60^{\circ}}{2}=30^{\circ}.$ $\boxed{\angle XAY=30^{\circ}}$
Appendix K Failure Cases of FRISM
In Fig. A4, we present a failure case of the proposed FRISM from the MathVista (Lu et al., 2024a) dataset. We compare the responses generated by the base VLM Qwen2.5-VL-32B-Instruct and the models merged with QwQ-32B through our FRISM. It can be seen that though showing significant performance improvements, merging VLMs with LRMs may introduce overthinking issues, which can lead to some failure cases.
Figure A4: A failure case of the proposed FRISM. The merged model first compute the correct answer before changing it to the wrong one due to overthinking.
@@
Despite the convincing results, the proposed FRISM suffers from the following limitations.
  • After merging with LRMs, the output length of VLMs can significantly increase, thus increasing the inference time and computation costs, which is a common problem within this field. Future studies may include jointly optimizing performance and efficiency.
  • Our reasoning objective uses injected subspace magnitude as a proxy because labeled VL reasoning data are unavailable in our setting. This choice is effective empirically but it is still a proxy rather than a direct supervision signal for reasoning transfer. Future studies may focus on the supervision signal for reasoning transfer during label-free training.
  • FRISM is fine-grained subspace-level merging based on posterior findings. The theoretical principles underlying this are worth further analysis. Also, as a common limitation of task vector-based merging methods, FRISM cannot be generalized to models with different sizes or structures.
Further transferring reasoning capabilities from reasoning language models to models with more modalities, model sizes, or even different structures is a significant future direction.